# A Unified Approach to Routing and Cascading for LLMs

### v2

@@ 1 Introduction
Lack of Deeper Understanding of Routing Algorithms
Further, the conditions under which current routing and cascading strategies are optimal, are not well understood. For routing, an extensive proof is required just to show that current strategies are close to optimal (Chen et al., 2022), while the theoretical analysis of current cascading algorithms is based on restrictive assumptions and does not provide optimality guarantees (Chen et al., 2023; Varshney & Baral, 2022). This lack of theoretical understanding hinders the development of more effective model selection strategies. More importantly, prior work fails to provide insights into the limitations of model selection strategies and cannot identify the conditions under which they are useful in practical scenarios. For instance, it is widely believed that one needs models of various sizes to benefit from cascading (Chen et al., 2023; Gupta et al., 2024; Khalili et al., 2022), but we show that this notion is incorrect.
@@ 2 Routing as Linear Optimization
In this section, we begin by defining a routing strategy as a function that maps queries to models. Next, we introduce the notation for quality and cost functions, demonstrating how the optimal routing strategy can be formulated to maximize a linear tradeoff between these two factors. Following this, we explain the process of selecting the best hyperparameters for the strategy. Lastly, we discuss the significance of quality estimation in routing and its impact on the effectiveness of the strategy.
@@
In routing, we seek to maximize the expected output quality of the selected model while adhering to a given cost budget $B$. Quality could measure model accuracy, user preference, or any other performance indicator. We define the quality function $q_{i}(x)$ as the output quality of model $m_{i}$ on query $x$, and the cost function $c_{i}(x)$ as the cost of running model $m_{i}$ on $x$.
@@
For a cost budget $B$, there exists a $\lambda\in\mathbb{R}^{+}$ and a $\gamma\in[0,1]$ such that the optimal routing strategy $s_{\textsc{opt}}$ equals $\gamma s_{\textsc{min}}^{\lambda}+(1-\gamma)s_{\textsc{max}}^{\lambda}$.
Theorem 1, continued. Furthermore, all routing strategies that have an expected cost that is exactly equal to $B$ and can be written as a convex combination of $s_{\textsc{min}}^{\lambda^{\prime}}$ and $s_{\textsc{max}}^{\lambda^{\prime}}$ for some $\lambda^{\prime}\in\mathbb{R}^{+}$ achieve the same optimal quality.
@@
Due to the second part of Theorem 1, we only need to find a set of hyperparameters $\lambda$ and $\gamma$ that achieve the cost budget. To determine these parameters, we estimate the cost of a strategy using a validation dataset $D$ that is representative of the query distribution $\mathcal{X}$. We then perform a hyperparameter search to find optimal values of $\lambda$ and $\gamma$. By leveraging several properties of routing strategies (see App. A), one can show that this hyperparameter search can be reduced to a single binary search over $\lambda$, enabling a quick and efficient hyperparameter optimization process.
@@ 3 Cascading as Sequential Routing
In this section, we reinterpret cascading as a sequence of routing problems. We then derive the optimal cascading strategy and explain how to determine the optimal hyperparameters by solving an additional optimization problem. Furthermore, we show how our approach improves upon the cascading strategies used in prior work. Finally, we examine the impact of post-hoc quality estimates on the effectiveness of cascading strategies.
@@
Notably, a model $m_{j}$ might perform poorly while $m_{j+1}$ performs exceptionally well on a given query. In such cases, the performance of $M_{1:j}$ will be worse than the performance of $M_{1:j-1}$, but $M_{1:j+1}$ could still provide a better outcome. Therefore, it is crucial to consider all supermodels $M_{1:j},\dots,M_{1:k}$ at step $j$ rather than making decisions solely based on immediate performance.
@@
The main difference between Theorem 2 and Theorem 1 is that not all combinations of hyperparameters $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}^{+}$ and $\gamma\in[0,1]$ that achieve cost budget $B$ are optimal. Instead, finding the optimal hyperparameters requires solving another optimization problem. Specifically, let us denote by $Q_{D}\colon(\mathbb{R}^{+})^{k+1}\rightarrow\mathbb{R}$, resp. $C_{D}\colon(\mathbb{R}^{+})^{k+1}\rightarrow\mathbb{R}$, the average quality, resp. cost, of a cascading strategy on a training dataset $D$ for a given set of hyperparameters $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}^{+}$ and $\gamma\in[0,1]$. Then, the optimal hyperparameters are the solution to the optimization problem
$\max_{\lambda_{1},\dots,\lambda_{k},\gamma}$ $Q_{D}(\lambda_{1},\dots,\lambda_{k},\gamma)$ (2)
s.t. $C_{D}(\lambda_{1},\dots,\lambda_{k},\gamma)\leqslant B.$
To solve this, we first assume $\lambda_{1}=\dots=\lambda_{k}=\lambda$ and apply the binary search technique from Section 2 to determine $\lambda$. This is then used as initialization for a hyperparameter optimization tool, Hyperopt [^2] https://github.com/hyperopt/hyperopt, to find the optimal values.
@@
Prior work on cascading has often relied on strong assumptions to simplify the strategy. The most common technique uses a threshold to decide whether to continue the cascade on an input $x$ (Chen et al., 2023; Gupta et al., 2024). Specifically, in step $j$, the cascade continues if $\hat{q}^{(j)}_{j-1}(x)<\tau_{j}$ for some $\tau_{j}\in\mathbb{R}$. Below, we outline the conditions under which this simplified approach is optimal.
@@ 4 Cascade Routing as Cascade Generalization
In this section, we first define cascade routing and explain how it generalizes both routing and cascading. We then derive the optimal cascade routing strategy and solve several of the additional challenges that arise when applying cascade routing in practice.
@@
While cascade routing extends cascading and can therefore use the same hyperparameter optimization scheme, it also introduces additional challenges which we address now.
@@
In cascading, the model order is predetermined, and the routing strategy only decides whether to proceed with the next model in the sequence. In contrast, cascade routing must dynamically determine the order in which models are computed. Despite this, both the estimated quality $\hat{q}^{(j)}_{M}(x)$ and cost $\hat{c}^{(j)}_{M}(x)$ of a supermodel $M$ are order-independent. Therefore, supermodels that contain the same models in a different order will have the same associated cost and quality. To mitigate this, we sort the models within the selected supermodel by cost and compute the cheapest one first. This approach aligns with cascading, where more expensive models are only used if cheaper models do not suffice.
@@
In cascading, the quality and cost must be computed for a maximum of $k$ supermodels at each step. However, in cascade routing, the number of supermodels grows exponentially, leading to the need to evaluate up to $2^{k}$ supermodels. This increase can become prohibitively costly, particularly since the model selection process must remain computationally negligible with respect to model computation. To mitigate this, we leverage so-called negative marginal gains. Specifically, if a model $m$ in a supermodel $M$ negatively impacts the quality-cost tradeoff, all supermodels containing all models in $M$ can be pruned from the search space. Since this negative contribution is quite common, this allows us to prune the search space significantly. More formally, this pruning operation relies on the following lemma:
@@ 5 Experimental Evaluation
We now evaluate the performance of cascade routing and demonstrate that it significantly outperforms all other strategies. Additionally, we show that our new cascading approach outperforms the threshold-based cascading method. For this purpose, we first conduct experiments on RouterBench (Hu et al., 2024), a benchmark specifically designed to evaluate routing and cascading (Section 5.1). Next, we test cascade routing on several additional benchmarks to evaluate its performance in more realistic scenarios (Section 5.2). In the appendix, we additionally perform an ablation study to examine the impact of various design choices in cascade routing on performance and runtime (App. F).
@@ 5.1 RouterBench
Similar to (Hu et al., 2024), we estimate quality and cost by adding zero-centered Gaussian noise to their true values. Both cost and quality estimates are modeled as linear functions fitted on these noisy signals. Thus, the quality estimate can be expressed as $\hat{q}_{W,b}(x)=W(q(x)+\epsilon)+b$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. A similar expression holds for the cost estimate. We define the variance of the noisy signal as $\sigma_{\text{ante}}^{2}$ before model computation (ex-ante estimates) and $\sigma_{\text{post}}^{2}$ after (post-hoc estimates). To explore different uncertainty levels, we vary the variances to simulate low-, medium-, and high-noise scenarios, with exact values for the variances given in Section D.1.
@@
((a)) Comparison of cascade routing with routing.
((b)) Comparison of cascade routing with cascading.
@@ 6 Related Work
There has also been specific research on quality estimation within cascading frameworks. Gupta et al. (2024) examine various measures of uncertainty in language model answers, evaluating their impact on cascading performance. Meanwhile, Jitkrittum et al. (2023) explore failure cases in cascading mechanisms that rely on uncertainty, introducing alternative quality measures that enhance cascade efficiency. Furthermore, Xue et al. (2023) apply cascading to majority voting for a single model to obtain a method called dynamic voting: the cascade stops depending on the aggregated answers of all previous model computations. Lastly, (Zhang et al., 2024a) propose the use of multi-objective quality metrics to guide cascading decisions and do not solely focus on accuracy.
@@ 7 Conclusion
In this work, we introduced a novel framework for routing and cascading that enabled us to propose theoretically optimal strategies for both paradigms. Further, we used this theoretical analysis to propose a new paradigm for model selection, cascade routing, which combines the benefits of both routing and cascading. We showed that cascade routing can significantly outperform its baselines, especially with good quality and cost estimates. We also find that our new cascading strategy significantly outperforms existing approaches to cascading, showing our theoretical analysis also leads to practical gains. Our work not only provides a theoretical foundation for model selection strategies but also opens up new avenues for future research in this area, especially in the direction of obtaining more accurate quality and cost estimates.
@@
Our work has the potential to significantly impact the field of model selection strategies. By providing a theoretical foundation for routing and cascading, we have shown that these strategies can be significantly improved by using more accurate quality and cost estimates. Cascade routing combines the strengths of both routing and cascading and offers a more flexible and effective model selection strategy. Furthermore, by underscoring the importance of quality estimation, we have highlighted a critical area for future research in model selection strategies that could lead to further improvements in this area.

### v3

@@ 1 Introduction
Lack of Deeper Understanding
Further, the conditions under which current routing and cascading strategies are optimal, are not well understood. For routing, an extensive proof is required just to show that current strategies are close to optimal (Chen et al., 2022), while the theoretical analysis of cascading does not provide optimality guarantees (Chen et al., 2023; Varshney & Baral, 2022). This lack of theoretical understanding hinders the development of more effective model selection strategies. Moreover, prior work fails to provide insights into the limitations of model selection strategies and cannot identify the conditions under which they are useful in practical scenarios. For instance, it is widely believed that one needs models of various sizes to benefit from cascading (Chen et al., 2023; Gupta et al., 2024; Khalili et al., 2022), but we show that this notion is incorrect.
@@ 2 Routing as Linear Optimization
In this section, we begin by defining a routing strategy as a function that maps queries to models. Next, we introduce the notation for quality and cost functions, demonstrating how the optimal routing strategy can be formulated to maximize a linear tradeoff between these two factors. Lastly, we give an illustrative example and discuss the significance of quality estimation in routing and its impact on the effectiveness of the strategy.
@@
In routing, we seek to maximize the output quality of the selected model while adhering to a given cost budget $B$. We define the quality function $q_{i}(x)$ as the output quality of model $m_{i}$ on query $x$, and the cost function $c_{i}(x)$ as the cost of running model $m_{i}$ on $x$. Quality could measure model accuracy, user preference, or any other performance indicator. Cost could measure either monetary costs or latency, depending on the use case.
@@
For a cost budget $B$, there exists a $\lambda\in\mathbb{R}^{+}$ and a $\gamma\in[0,1]$ such that the optimal routing strategy $s_{\textsc{opt}}$ equals $\gamma s_{\textsc{min}}^{\lambda}+(1-\gamma)s_{\textsc{max}}^{\lambda}$. Theorem 1, continued. Furthermore, all routing strategies that have an expected cost that is exactly equal to $B$ and can be written as a convex combination of $s_{\textsc{min}}^{\lambda^{\prime}}$ and $s_{\textsc{max}}^{\lambda^{\prime}}$ for some $\lambda^{\prime}\in\mathbb{R}^{+}$ achieve the same optimal quality.
In App. A, we show how to obtain the optimal $\lambda$ and $\gamma$ for a cost budget $B$ using a validation dataset $D$. In Algorithm 1, we provide pseudocode for the optimal routing algorithm.
Algorithm 1 Optimal Routing Algorithm
Input: input query $x$, quality estimator $\hat{q}_{i}$, cost estimator $\hat{c}_{i}$, tradeoff parameters $\lambda$ and $\gamma$
Output: Model index $i$ to be used for query $x$
1:  $\tau_{i}(x,\lambda):=\hat{q}_{i}(x)-\lambda\hat{c}_{i}(x)$
2:  $\tau_{\text{max}}(x,\lambda):=\max_{i\in\{1,\dots,k\}}\tau_{i}(x,\lambda)$
3:  $\textrm{best}:=\{i\in\{1,\dots,k\}|\tau_{i}(x,\lambda)=\tau_{\text{max}}(x,\lambda)\}$
[…]
7:     return min_cost_index
8:  else
9:     return max_cost_index
10:  end if
@@
Example
To illustrate routing, consider a scenario with two models $m_{1}$ and $m_{2}$. The estimated cost $\hat{c}(x)=(0.9,1)$ is constant across queries and slightly lower for $m_{1}$ than for $m_{2}$. The quality is estimated based on the category of the query, which is either math, code, or generic. For instance, let $\hat{q}(x_{\textrm{math}})=(0.8,0.5)$, $\hat{q}(x_{\textrm{code}})=(0.5,0.8)$, and $\hat{q}(x_{\textrm{generic}})=(0.8,0.9)$. For $\lambda=1$ and $\gamma=0.7$, the cost-quality tradeoff is highest for $m_{1}$, resp. $m_{2}$, on math, resp. code, queries, and is equal for both models on generic queries. Thus, the router will select $m_{1}$, resp. $m_{2}$, for math, resp. code, queries, and select $m_{1}$ with a probability of $0.7$ and $m_{2}$ with a probability of $0.3$ for generic queries.
@@ 3 Cascading as Sequential Routing
In this section, we reinterpret cascading as a sequence of routing problems. Furthermore, we show how our approach improves upon the cascading strategies used in prior work. Finally, we examine the impact of post-hoc quality estimates on the effectiveness of cascading strategies.
@@
Notably, while the action associated with supermodels $M_{1:j},\dots,M_{1:k}$ is the same, namely continuing the cascade, it is important to consider all these supermodels. Indeed, a model $m_{j}$ might perform poorly while $m_{j+1}$ performs exceptionally well on a given query. In such cases, the quality-cost tradeoff of $M_{1:j}$ will be worse than the tradeoff of $M_{1:j-1}$, but $M_{1:j+1}$ could still provide a better outcome. Therefore, it is crucial to consider all supermodels $M_{1:j},\dots,M_{1:k}$ at step $j$ rather than making decisions solely based on immediate performance.
@@
In App. B, we explain how to obtain the hyperparameters $\lambda_{1},\dots,\lambda_{k}$ and $\gamma$ for a given cost budget $B$ using a validation dataset $D$. In Algorithm 2, we provide pseudocode for the optimal cascading algorithm, illustrating the steps involved in selecting the best model for a given query.
Algorithm 2 Optimal Cascading Algorithm
Input: input query $x$, current step $j$, quality estimator $\hat{q}_{i}^{(j)}$, cost estimator $\hat{c}_{i}^{(j)}$, tradeoff parameters $\lambda_{j}$ and $\gamma$
Output: Whether to stop the cascade or continue with the next model $m_{j}$
1:  for $i=j-1$ to $k$ do
2:     $\hat{q}^{(j)}_{1:i}(x):=\mathbb{E}_{\hat{q}^{(j)}}[\max(\hat{q}_{1}^{(j)}(x),\dots,\hat{q}_{i}^{(j)}(x))]$
3:     $\hat{c}^{(j)}_{1:i}(x):=\sum_{l=1}^{i}\hat{c}_{l}^{(j)}(x)$
4:  end for
5:  $\textrm{index}:=\textrm{Router}\Big{(}x,(\hat{q}^{(j)}_{1:j-1}(x),\dots,\hat{q}^{(j)}_{1:k}(x)),$
[…]
8:     return stop
9:  else
10:     return run $\textrm{run }m_{j}$
11:  end if
Example
To illustrate the optimal cascading strategy, consider once again a scenario with two models $m_{1},m_{2}$ with costs $\hat{c}^{(1)}(x)=(0.5,1)$ and $\hat{q}^{(1)}(x)=(0.5,0.8)$ (ignoring uncertainty). In a cascade, we will always run $m_{1}$ for a query $x$. Thus, we obtain the first model’s output. Based on this output, we can adjust the quality and cost estimates. Suppose, for instance, that the model output is very long and its confidence in its own answer is only $30\%$. Then we update, for instance, $\hat{q}^{(2)}(x)=(0.3,0.6)$ and $\hat{c}^{(2)}(x)=(1,2)$. For $\lambda_{2}=1$, we would now stop the cascade and return the output of $m_{1}$ since the cost-quality tradeoff is highest for the supermodel $\{m_{1}\}$. If, instead, $\lambda_{2}=0.1$, we would run $m_{2}$ since the cost-quality tradeoff is highest for the supermodel $\{m_{1},m_{2}\}$. In this case, the cascade would return the output of $m_{2}$.
@@
Prior work on cascading has often relied on strong assumptions to simplify the strategy, using a treshold-strategy as an approximation of the optimal cascade. Specifically, in step $j$, the cascade continues if $\hat{q}^{(j)}_{j-1}(x)<\tau_{j}$ for some threshold $\tau_{j}\in\mathbb{R}$. To the best of our knowledge, all existing works can be seen as a specific instantiation of this thresholding scheme with cost and quality estimators that depend on the specific application used (Chen et al., 2023; Damani et al., 2024; Gupta et al., 2024; Jitkrittum et al., 2023b; Nie et al., 2024). Below, we outline the conditions under which this simplified approach is optimal.
@@ 4 Cascade Routing as Cascade Generalization
In this section, we first define cascade routing and explain how it generalizes both routing and cascading. We then derive the optimal cascade routing strategy and solve several of the additional challenges that arise when applying cascade routing in practice. Finally, we provide an illustrative example.
@@
In Algorithm 3, we provide pseudocode for the cascade routing algorithm. While cascade routing is a seemingly simple extension of cascading, it also introduces additional challenges which we address now.
Algorithm 3 Optimal Cascade Routing Algorithm
Input: input query $x$, model indices run so far $\{i_{1},...,i_{j-1}\}$, quality estimator $\hat{q}_{i}^{(j)}$, cost estimator $\hat{c}_{i}^{(j)}$, tradeoff parameters $\lambda_{j}$ and $\gamma$
Output: Index of the next model to run
1:  $\mathcal{S}=[M\subset\{1,\dots k\}|\forall l\in\{1,\dots,j-1\}:i_{l}\in M]$
2:  for $M\in\mathcal{S}$ do
3:     $\hat{q}^{(j)}_{M}(x):=\mathbb{E}_{\hat{q}^{(j)}}[\max_{l\in M}(\hat{q}_{l}^{(j)}(x))]$
4:     $\hat{c}^{(j)}_{M}(x):=\sum_{l\in M}\hat{c}_{l}^{(j)}(x)$
5:  end for
6:  $\textrm{index}:=\textrm{Router}\Big{(}x,(\hat{q}^{(j)}_{M}(x)\text{ for }M\in\mathcal{S}),$
[…]
8:  $\textrm{possibilities}:=\mathcal{S}_{\textrm{index}}\setminus\{i_{1},...,i_{j-1}\}$
9:  if $\textrm{possibilities}=\emptyset$ then
10:     return stop
11:  else
[…]
13:     return min_cost_index
14:  end if
@@
In cascading, the model order is predetermined, and the routing strategy only decides whether to proceed with the next model in the sequence. In contrast, cascade routing must dynamically determine the order in which models are computed. Despite this, both the estimated quality $\hat{q}^{(j)}_{M}(x)$ and cost $\hat{c}^{(j)}_{M}(x)$ of a supermodel $M$ are order-independent. Therefore, supermodels that contain the same models in a different order will have the same cost and quality. To mitigate this, we sort the models within the selected supermodel by cost and compute the cheapest one first (illustrated in Lines 12-13 of Algorithm 3). This approach aligns with cascading, where more expensive models are only used if cheaper models do not suffice.
@@
In cascading, the quality and cost must be computed for a maximum of $k$ supermodels at each step. However, in cascade routing, the number of supermodels grows exponentially, leading to the need to evaluate up to $2^{k}$ supermodels. This increase can become prohibitively costly, particularly since the model selection process must remain computationally negligible with respect to model computation. To mitigate this, we leverage so-called negative marginal gains. It can be shown (see App. C) that if a model $m$ in a supermodel $M$ negatively impacts the quality-cost tradeoff, all supermodels containing all models in $M$ can be pruned from the search space. For example, if $m_{1}$ negatively affects the quality-cost tradeoff of the supermodel $\{m_{1},m_{2}\}$, we can prune all supermodels that contain both $m_{1}$ and $m_{2}$. Since this negative contribution is quite common, this allows us to prune the search space significantly. More formally, this pruning operation relies on the following lemma:
@@
Example
To illustrate the optimal cascade routing strategy, consider a scenario with two models $m_{1}$ and $m_{2}$ with costs $\hat{c}^{(1)}(x)=(0.5,1)$ and $\hat{q}^{(1)}(x)=(0.5,0.8)$. In contrast to cascading, we do not necessarily need to run $m_{1}$ first. In this scenario, if $\lambda_{1}=0.1$, we would immediately run $m_{2}$. While we would most likely stop after $m_{2}$, it is possible that its output is so bad that we update the quality estimator to $\hat{q}^{(2)}(x)=(0.5,0.1)$. In this case, for $\lambda_{2}=0.1$, we would run $m_{1}$ next and return its output. If $\lambda_{1}=1$, we come into the more classical cascading scenario explained before where $m_{1}$ is run first.
@@
((a)) Comparison of cascade routing with routing.
((b)) Comparison of cascade routing with cascading.
@@ 5 Experimental Evaluation
We now evaluate the performance of cascade routing and demonstrate that it significantly outperforms all other strategies. Additionally, we show that our new cascading approach outperforms the threshold-based cascading method. For this purpose, we first conduct experiments on RouterBench (Hu et al., 2024), a benchmark specifically designed to evaluate routing and cascading (Section 5.1). Next, we test cascade routing on several additional benchmarks to evaluate its performance in more realistic scenarios (Section 5.2). In the appendix, we perform an ablation study to examine the impact of various design choices in cascade routing on performance and runtime (App. F). Finally, in App. H, we show detailed results as well as cost-quality tradeoff curves for several benchmarks.
@@ 5.1 RouterBench
Similar to (Hu et al., 2024), we estimate quality and cost by adding zero-centered Gaussian noise to their true values. Both cost and quality estimates are modeled as linear functions fitted on these noisy signals. Thus, the quality estimate can be expressed as $\hat{q}_{W,b}(x)=W(q(x)+\epsilon)+b$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. A similar expression holds for the cost estimate. We define the variance of the noisy signal as $\sigma_{\text{ante}}^{2}$ before model computation (ex-ante estimates) and $\sigma_{\text{post}}^{2}$ after (post-hoc estimates).
To explore different uncertainty levels, we vary the variances to simulate low-, medium-, and high-noise scenarios, with exact values for the variances given in Section D.1.
@@ Results
We highlight that the cost estimator for SWE-Bench is latency-based, computing cost as the time it takes to complete the task. In contrast, the estimator for Minerva Math and LiveCodeBench uses the cost of the generation. Thus, cascade routing can adapt to different cost estimators.
@@ 6 Related Work
There has also been specific research on quality estimation within cascading frameworks. Gupta et al. (2024) examine various measures of uncertainty in language model answers, evaluating their impact on cascading performance. Meanwhile, Jitkrittum et al. (2023a) explore failure cases in cascading mechanisms that rely on uncertainty, introducing alternative quality measures that enhance cascade efficiency. Furthermore, Xue et al. (2023) apply cascading to majority voting for a single model to obtain a method called dynamic voting: the cascade stops depending on the aggregated answers of all previous model computations. Lastly, (Zhang et al., 2024a) propose the use of multi-objective quality metrics to guide cascading decisions and do not solely focus on accuracy.
All works mentioned here can be seen as an instantiation of the thresholding mechanism outlined in Corollary 1 with application-specific quality and cost estimates.
@@ 7 Conclusion
In this work, we introduced a novel framework for routing and cascading that enabled us to propose theoretically optimal strategies for both paradigms. Further, we used this analysis to propose a new paradigm for model selection, cascade routing, which combines the benefits of routing and cascading. We showed that cascade routing can significantly outperform its baselines, especially with good quality and cost estimates. We also find that our new cascading strategy significantly outperforms existing approaches to cascading, showing our theoretical analysis also leads to practical gains.
@@
Our work can significantly impact the field of model selection strategies. By providing a theoretical foundation for routing and cascading, we have shown that these strategies can be improved by using more accurate quality and cost estimates. Cascade routing combines the strengths of both routing and cascading and offers a more flexible and effective model selection strategy. Furthermore, by underscoring the importance of quality estimation, we highlighted a critical area for future research in model selection strategies that could lead to further improvements in this area.